OpenStudy (anonymous):

sin^2(-theta) + cos^2(-theta) = 1 Establish the identity.

OpenStudy (anonymous):

if a be the base, b the opposite side and c the hypotenuse sin t = b/c cos t= a/c sin^2t + cos^2t= (a^2 + b^2 ) /c^2 by pythagoras thm a^2 + b^2 = c^2 hence proved

OpenStudy (anonymous):

Thanks man!